# Transform and Entropy Coding in AV2

### v1

@@ Abstract
AV2 is the successor to the AV1 royalty-free video coding standard developed by the Alliance for Open Media (AOMedia). Its primary objective is to deliver substantial compression gains and subjective quality improvements while maintaining low-complexity encoder and decoder operations. This paper describes the transform, quantization and entropy coding design in AV2, including redesigned transform kernels and data-driven transforms, expanded transform partitioning, and a mode & coefficient dependent transform signaling. AV2 introduces several new coding tools including Intra/Inter Secondary Transforms (IST), Trellis Coded Quantization (TCQ), Adaptive Transform Coding (ATC), Probability Adaptation Rate Adjustment (PARA), Forward Skip Coding (FSC), Cross Chroma Component Transforms (CCTX), Parity Hiding (PH) tools and improved lossless coding. These advances enable AV2 to deliver the highest quality video experience for video applications at a significantly reduced bitrate.
@@ I Introduction
Global demand for digital video has surged over the last decade, making it the dominant driver of internet bandwidth consumption and a major source of compute and storage load in content delivery systems [ericsson-2025, itu-usage, cisco-annual]. This demand is fueled by applications such as video streaming and transcoding, video-on-demand (VOD), real-time communications (RTC), and emerging domains including immersive AR/VR, AI-generated content, and cloud gaming [kumar2024cloud, bukhari2023video, meta-reels]. The widespread adoption of 4K and 8K displays, higher frame rates, high dynamic range (HDR), and volumetric or interactive content is further expected to increase the demands on video compression technology [chiariotti2021survey]. These trends place a significant strain on delivery infrastructures, where bandwidth and storage remain costly resources.
The Alliance for Open Media (AOMedia) has developed the AV2 video codec as the successor to AV1 building on the same royalty-free foundation [aom]. Released in 2018, AV1 delivered 30% bitrate savings over VP9 at equivalent quality and has since been deployed widely across web platforms and devices [han2021av1]. AV2 extends this trajectory by providing an additional 30% bitrate reduction over AV1. The design of AV2 emphasizes reduced implementation and hardware cost ensuring practicality for both software and hardware deployments.
@@
On the entropy coding side, AV2 retains the multi-symbol arithmetic coding (MS-AC) engine of AV1 but improves it’s efficiency through several new tools. The probability adaptation rate adjustment (PARA) tool provides an optimized probability adaptation for MS-AC engine by adjusting the adaptation rate per syntax. Adaptive transform coding (ATC) improves the coefficient context modeling and unifies coefficient scanning rules over AV1. Adaptive truncated Rice coding enhances the representation of large coefficient magnitudes, trellis-coded quantization (TCQ) introduces an RD-optimized framework for coefficient coding, and forward skip coding (FSC) enables compact representation of directly coded residuals enhancing coding efficiency for screen-content and denser residuals. Lastly, a new parity hiding (PH) tool reduces signaling overhead associated with the DC coefficient magnitudes as an alternative to TCQ.
These coding tools provide significant efficiency gains across both natural video and screen content while preserving high throughput and low-latency operation. This paper provides an overview of the transform and entropy coding design in AV2. Section II reviews the AV1 framework to establish the legacy AV1 design. Section III details the transform and entropy coding coding methods of AV2. Section IV summarizes compression performance per coding tool.
@@ II-A Transform Block Partitioning in AV1
AV1 supports TB sizes ranging from $4\times 4$ up to $64\times 64$, including both square and rectangular configurations such as $\times N/2, N×N$$\times N/4$, $N/2\times N$, and $N/4\times N$ [han2021av1]. Transform partitioning is applied recursively to split a CB into multiple TBs.
@@ II-B Primary Transforms in AV1
The DCT transform is effective for compacting smoother residuals and homogeneous regions, whereas ADST and FLIPADST transforms provide better energy compaction for directional edges often encountered in inter predicted residuals. The IDTX is useful for preserving sharp discontinuities such as those encountered in screen content [chen2019av1tools, han2021av1].
@@ II-E Entropy Coding Engine
AV1 uses a context-based multi-symbol arithmetic coder (MS-AC) to efficiently encode syntax elements in the bitstream [han2021av1]. Unlike binary arithmetic coders (BAC) which can represent only a single binary outcome per coding step [langdon1984introduction, said2003arithmetic], the MS-AC engine in AV1 supports up to 16 symbols per syntax element with probabilities represented using cumulative distribution function (CDF) tables. MS-AC improves coding efficiency for syntax elements with incremental values by avoiding the explicit binarization traditionally required in BACs.
@@ II-F Coefficient Level Map and Sign Coding
The transform compacts residuals into a smaller set of transform coefficients. These coefficients are then quantized and coded with a level-map representation.
The coefficient coding process in AV1 begins by checking whether the TB contains any nonzero coefficients. If all coefficients are quantized to zero, an all_zero syntax element is signaled with a value of ‘1’ as the first element of the TB-level syntax which terminates future coding of coefficient related syntax. If the block contains at least one nonzero coefficient all_zero sends a value of ‘0’, and an end-of-block (EOB) syntax element is coded in the bitstream. The EOB syntax is transmitted using a hierarchical scheme that identifies a coarse group of coefficient positions followed by finer bins to determine the exact location of the last significant coefficient.
Once the EOB syntax is coded transform coefficients are signaled in a pre-determined scan order. AV1 uses several coefficient scan orders depending on transform type and size: zig-zag, up-right diagonal, bottom-left diagonal, row, or column. One-dimensional transform types (horizontal or vertical) use row or column scans, while two-dimensional transforms use diagonal or zig-zag orders.
@@
  • Base Range (BR): Uses a 4-symbol syntax element that uses contexts for entropy modeling dependent on block size and the level values of neighboring coefficients. This range covers level values [0, 1, 2] with an additional symbol indicating magnitudes larger than $\geq$2 which determines whether an additional pass is need if the magnitude is greater. This pass uses a reverse scan order.
  • Low Range (LR): Coded with another 4-symbol syntax element that loops 4 times depending on how large the actual level is. This range covers level values in the range [3, 4, … 14] and can also indicate if level value is $\geq$ 14. LR pass uses a reverse scan order.
  • High Range (HR): This range covers largest level values $\geq 14$ by using exponential-Golomb codes. The difference between the actual level value and 15 is computed since the previous pass signals $\geq$ 14 which means the value is at least 15. This passes use a forward scan order different than the BR and LR passes.
Coefficient signs are signaled as bypass-coded bits in the HR pass, with only the DC coefficient assigned a sign context.
@@ III-B Primary Transforms and AV2 Improvements
AV2 introduces data-driven transforms (DDTs) for $8\times 8$ and $16\times 16$ inter blocks. These kernels are trained offline via Karhunen–Loève Transformation (KLT) on prediction residuals. Each DDT includes a learned kernel and its flipped variant (e.g., DDT8, FLIPDDT8). DDTs use the same transform type syntax as FLIPADST/ADST and integrate cleanly into the transform pipeline e.g. the are used in-place of ADST variants shown in Table I reusing the same syntax. For inter blocks, the 8-point DDT replaces the learned ADST (L-ADST), and the 16-point DDT replaces the fixed DST-7 kernel. Directional transforms DST-7 and L-ADST still remain available for intra-coded blocks.
@@ III-D Intra/Inter Secondary Transforms
Separable spatial transforms constitute a fundamental tool for decorrelating prediction residuals in major video coding standards, including H.264/AVC [Wieg2003], HEVC [sull2012], AV1 [han2021av1], and VVC [xin2021]. However, their effectiveness is limited for residual signals containing arbitrarily oriented edges and complex textures commonly observed in natural video content. Non-separable transforms, while capable of better modeling such patterns, incur significantly higher computational complexity [Arru2014, Taka2013]. To mitigate this cost, non-separable secondary transforms first introduced in [Xin2016], are applied as an intermediate stage between the primary transform and quantization. Further refinements, including restricting the transform to low-frequency coefficients, zeroing out high-frequency components, and limiting kernel sizes [Koo2019], have made these approaches practical for modern codecs. In AV2, Intra/Inter Secondary Transforms (IST) are introduced as part of the AV2 transform coding framework to improve residual energy compaction while maintaining manageable complexity.
@@ III-E Cross-chroma Component Transforms (CCTX)
The well-known YCbCr color space provides good perceptual uniformity and overall compression efficiency. However, local correlations can still exist among the three YCbCr planes. While existing tools such as chroma from luma (CfL) [trudeau2018predicting] effectively reduce the luma-chroma correlation, a new tool called cross-chroma component transforms (CCTX) has been introduced in AV2 to specifically address the correlation between the two chroma planes (CbCr) after a primary transforms.
@@ III-H Trellis Coded Quantization
In AV2, a Trellis Coded Quantization (TCQ) mode is introduced as an alternative to the scalar quantization. TCQ can be enabled at the sequence or frame level, allowing switching between the TCQ method and the scalar or Parity Hiding (PH) methods on a frame-by-frame basis. When TCQ mode is enabled for a frame, the TCQ method is applied to luma TBs while the scalar quantization method is still used for TBs that use 1D scan, chroma blocks, and blocks that use Forward Skip Coding (FSC).
Similar to earlier TCQ methods, such as that described in the JPEG 2000 extensions [t.801], the AV2 TCQ mode defines two quantizers, labeled as Q0 and Q1. As shown in Figure 5, the non-zero reconstruction levels of Q1 are offset by a one-half step relative to Q0. The selection of these quantizers is governed by a finite state machine as shown in Figure 6. The initial state is reset to zero at the start of decoding each TB. After each quantized coefficient value is determined, the current state transitions to one of two possible next-states, based on the parity (least-significant bit) of the current coded coefficient, which influences the future states of the next coefficients.
@@
III-H1 TCQ Low-level Syntax
@@
III-H2 TCQ Encoder Algorithm
The reference encoder performs coefficient optimization using the Viterbi dynamic programming algorithm, which jointly optimizes both the quantization decisions and the end-of-block (EOB) position. The encoder maintains nine states: one for all-zero coefficients, and eight coded states corresponding the states defined in the TCQ state diagram in Figure  6. The algorithm is executed in two passes.
In the first pass, the algorithm proceeds from the last non-zero coded coefficient (i.e. identified through simple quantization) and iterates through each coefficient position in reverse diagonal scanning order. At each position, quantization decisions (represented by the transition arrows in Figure  8) are evaluated to update the rate-distortion costs for each potential next-state. When multiple decisions lead to the same next-state, only the decision with the lowest rate-distortion cost is retained. The iterative process continues until the final coefficient position is reached.
@@ III-I Probability Adaptation Rate Adjustment (PARA)
In AV2, separate adjustment parameters (denoted as $\mathbf{a}$ in Figure  9) are jointly optimized per-context (i.e., group of symbols associated with each row of the CDF table), and different time intervals defined by the $r_{C}(n)$ in Equation 13 so that each time the interval $\mathcal{T}$ is defined as:
@@ III-J1 Adaptive Coefficient Coding
ATC modifies the coefficient coding syntax in each region for AV2. In the LF region, the BR level range is extended to to cover the level values $\text{BR}_{\text{LF}}=[0,1,2,3,4,\geq 4)$ using a 6-symbol alphabet. The final symbol $\geq 4$ serves as an escape character indicating that the actual value may exceed 4 which may trigger another coding pass. In the Default region, the BR range remains identical to AV1: $\text{BR}_{\text{Default}}=[0,1,2,\geq 2)$ including level values $0,1,2$ and the escape character for greater $\geq 2$.
@@
Default Any $\min(nstats,\ 3)+\text{V}_{\text{offset}}$ 0–7
@@
For the LR coding pass the context index is derived similarly from nstats with rules varying by channel (luma or chroma) and region (LF or Default) as detailed in Table VII. Luma uses a clip of 6, while chroma uses a clip of 3. Additional offsets are applied based on coefficient position (e.g., DC vs. non-DC) in the LF region.

### v2

@@ Abstract
AV2 is the successor to the AV1 video coding format developed by the Alliance for Open Media (AOMedia). Its primary objective is to deliver substantial compression gains and subjective quality improvements while maintaining low-complexity encoder and decoder operations. This paper describes the transform, quantization and entropy coding design in AV2, including redesigned transform kernels and data-driven transforms, expanded transform partitioning, and a mode & coefficient dependent transform signaling. AV2 introduces several new coding tools including Intra/Inter Secondary Transforms (IST), Trellis Coded Quantization (TCQ), Adaptive Transform Coding (ATC), Probability Adaptation Rate Adjustment (PARA), Forward Skip Coding (FSC), Cross Chroma Component Transforms (CCTX), Parity Hiding (PH) tools and improved lossless coding. These advances enable AV2 to deliver the highest quality video experience for video applications at a significantly reduced bitrate.
@@ I Introduction
Global demand for digital video has surged over the last decade, making it the dominant driver of internet bandwidth consumption and a major source of compute and storage load in content delivery systems [9, 12, 8]. This demand is fueled by applications such as video streaming and transcoding, video-on-demand (VOD), real-time communications (RTC), and emerging domains including immersive AR/VR, AI-generated content, and cloud gaming [14, 5, 16]. The widespread adoption of 4K and 8K displays, higher frame rates, high dynamic range (HDR), and volumetric or interactive content is further expected to increase the demands on video compression technology [7]. These trends place a significant strain on delivery infrastructures, where bandwidth and storage remain costly resources.
The Alliance for Open Media (AOMedia) has developed the AV2 video codec as the successor to AV1 [10]. Released in 2018, AV1 delivered 30% bitrate savings over VP9 at equivalent quality and has since been deployed widely across web platforms and devices [11]. AV2 extends this trajectory by providing an additional 30% bitrate reduction over AV1. The design of AV2 emphasizes reduced implementation and hardware cost ensuring practicality for both software and hardware deployments.
@@
In terms of quantization and coefficient coding, AV2 improves the design of the default scalar quantizer from AV1 and builds on top it to enable trellis-coded quantization (TCQ) to achieve higher compression efficiency with modest additional complexity. AV2 retains the multi-symbol arithmetic coding (MS-AC) engine of AV1 and further improves its efficiency through several new entropy coding tools. The probability adaptation rate adjustment (PARA) tool provides an optimized probability adaptation for MS-AC engine by adjusting the adaptation rate per syntax. Adaptive transform coding (ATC) improves the coefficient context modeling and unifies coefficient scanning rules over AV1. Adaptive truncated Rice coding enhances the representation of large coefficient magnitudes, and forward skip coding (FSC) enables compact representation of directly coded residuals enhancing coding efficiency for screen-content and denser residuals. Lastly, a new parity hiding (PH) tool reduces signaling overhead associated with the DC coefficient magnitudes, and can bring coding gains in the absence of TCQ.
These coding tools provide significant efficiency gains across both natural video and screen content while preserving high throughput and low-latency operation. This paper provides an overview of the transform and entropy coding design in AV2. Section II reviews the AV1 framework to establish the legacy AV1 design. Section III details the transform and entropy coding methods of AV2. Section IV summarizes compression performance per coding tool.
@@ II-A Transform Block Partitioning in AV1
AV1 supports TB sizes ranging from $4\times 4$ up to $64\times 64$, including both square and rectangular configurations such as $\times N/2, N×N$$\times N/4$, $N/2\times N$, and $N/4\times N$ [11]. Transform partitioning is applied recursively to split a CB into multiple TBs.
@@ II-B Primary Transforms in AV1
The DCT transform is effective for compacting smoother residuals and homogeneous regions, whereas ADST and FLIPADST transforms provide better energy compaction for directional edges often encountered in inter predicted residuals. The IDTX is useful for preserving sharp discontinuities such as those encountered in screen content [6, 11].
@@ II-E Entropy Coding Engine
AV1 uses a context-based multi-symbol arithmetic coder (MS-AC) to efficiently encode syntax elements in the bitstream [11]. Unlike binary arithmetic coders (BAC) which can represent only a single binary outcome per coding step [15, 18], the MS-AC engine in AV1 supports up to 16 symbols per syntax element with probabilities represented using cumulative distribution function (CDF) tables. MS-AC improves coding efficiency for syntax elements with incremental values by avoiding the explicit binarization traditionally required in BACs.
@@ II-F Coefficient Level Map and Sign Coding
The transform compacts residuals into a smaller set of transform coefficients. These coefficients are then quantized and coded with a level-map representation, where the level denotes coefficient magnitude.
The coefficient coding process in AV1 begins by checking whether the TB contains any nonzero coefficients. If all coefficients are quantized to zero, an all_zero syntax element is signaled with a value of ‘1’ as the first element of the TB-level syntax which terminates future coding of coefficient related syntax. If the block contains at least one nonzero coefficient all_zero sends a value of ‘0’, and an end-of-block (EOB) syntax element is coded in the bitstream. The EOB syntax is transmitted using a hierarchical scheme that identifies a coarse group of coefficient positions followed by finer bins to determine the exact location of the last significant coefficient. The number of bits required for EOB signaling depends on residual energy compaction with fewer bits spent when coefficients are clustered close towards the DC term.
Once the EOB syntax is coded, transform coefficients are signaled in a pre-determined scan order. AV1 uses several coefficient scan orders depending on transform type and size: zig-zag, up-right diagonal, bottom-left diagonal, row, or column. One-dimensional transform types (horizontal or vertical) use row or column scans, while two-dimensional transforms use diagonal or zig-zag orders, as illustrated in Fig. 2.
Figure 2: AV1 scan orders and context neighborhoods (green) for coefficient coding; numbers show forward scan rank, and the orange cell is the current coefficient to be coded.
@@
  • Base Range (BR): A 4-symbol syntax element that is context-coded using neighboring decoded levels (Fig. 2). This range covers level values $\{0,1,2\}$ with an additional escape symbol indicating magnitudes $\geq 3$ which determines whether an additional pass is needed. This pass uses a reverse scan order.
  • Low Range (LR): A 4-symbol syntax element that may iterate up to four times, extending the representation to levels $3,\ldots,14]$ with an escape for level $\geq 15$. LR uses a reduced context neighborhood (asterisks in Fig. 2) and reverse scan order.
  • High Range (HR): For magnitudes $\geq 15$, the residual magnitude beyond 15 is bypass-coded using exponential-Golomb codes. This pass uses a forward scan order different than the BR and LR passes.
Coefficient signs are bypass-coded in the HR stage, except that the DC coefficient uses a dedicated sign context.
@@ III-B Primary Transforms and AV2 Improvements
AV2 introduces data-driven transforms (DDTs) for $8\times 8$ and $16\times 16$ inter blocks. These kernels are trained offline via Karhunen–Loève Transformation (KLT) on prediction residuals. Each DDT includes a learned kernel and its flipped variant (e.g., DDT8, FLIPDDT8). DDTs use the same transform type syntax as FLIPADST/ADST and integrate cleanly into the transform pipeline, e.g., they are used in-place of ADST variants shown in Table I reusing the same syntax. For inter blocks, the 8-point DDT replaces the learned ADST (L-ADST), and the 16-point DDT replaces the fixed DST-7 kernel. Directional transforms DST-7 and L-ADST still remain available for intra-coded blocks.
@@ III-D Intra/Inter Secondary Transforms
Separable spatial transforms constitute a fundamental tool for decorrelating prediction residuals in major video coding standards/formats, including H.264/AVC [22], HEVC [19], AV1 [11], and VVC [24]. However, their effectiveness is limited for residual signals containing arbitrarily oriented edges and complex textures commonly observed in natural video content. Non-separable transforms, while capable of better modeling such patterns, incur significantly higher computational complexity [3, 20]. To mitigate this cost, non-separable secondary transforms first introduced in [23], are applied as an intermediate stage between the primary transform and quantization. Further refinements, including restricting the transform to low-frequency coefficients, zeroing out high-frequency components, and limiting kernel sizes [13], have made these approaches practical for modern codecs. In AV2, Intra/Inter Secondary Transforms (IST) are introduced as part of the AV2 transform coding framework to improve residual energy compaction while maintaining manageable complexity.
@@ III-E Cross-chroma Component Transforms (CCTX)
The well-known YCbCr color space provides good perceptual uniformity and overall compression efficiency. However, local correlations can still exist among the three YCbCr planes. While existing tools such as chroma from luma (CfL) [21] effectively reduce the luma-chroma correlation, a new tool called cross-chroma component transforms (CCTX) has been introduced in AV2 to specifically address the correlation between the two chroma planes (CbCr) after a primary transforms.
@@ III-H Trellis Coded Quantization
Trellis coded quantization (TCQ) is a well-studied advanced quantization method that can improve coding efficiency with a modest increase in complexity [17]. AV2 supports flexible signaling of quantization methods – between the default scalar quantization (SQ), TCQ, and Parity Hiding (PH) methods – at both the sequence and frame levels. When TCQ is enabled for a frame, it is applied to luma transform blocks (TBs) with 2D scan, whereas TBs with 1D scan, chroma blocks, and blocks with Forward Skip Coding (FSC) all fall back to SQ.
Similar to existing methods – such as the version adopted by JPEG 2000 extensions [1] – the TCQ mode in AV2 defines two underlying sub-quantizers, labeled as Q0 and Q1. As Fig. 6 illustrates, the non-zero reconstruction levels of Q1 are offset by a one-half step relative to Q0. The sub-quantizer selection is governed by a finite-state machine where each state is associated with either Q0 or Q1 (see Fig. 7).
The initial state is reset to zero at the start of decoding each TB. As coefficients are processed in scan order, their magnitudes and parities are decoded using novel syntax that ensures consistent parity across multiple syntax passes. The parity bit of the decoded coefficient (i.e., even or odd) dictates the transition to one of two possible next states, which in turn determines which sub-quantizer will be applied for the next coefficient.
@@
III-H1 Syntax Adjustments for Coefficient Coding
@@
III-H2 Encoder Algorithm
The reference encoder performs coefficient optimization using the Viterbi dynamic programming algorithm, which jointly optimizes both the quantization decisions and the end-of-block (EOB) position. The encoder maintains nine states: one for all-zero coefficients, and eight coded states corresponding the states defined in the TCQ state diagram in Figure  7. The algorithm is executed in two passes.
In the first pass, the algorithm proceeds from the last non-zero coded coefficient (i.e. identified through simple quantization) and iterates through each coefficient position in reverse diagonal scanning order. At each position, quantization decisions (represented by the transition arrows in Figure  9) are evaluated to update the rate-distortion costs for each potential next-state. When multiple decisions lead to the same next-state, only the decision with the lowest rate-distortion cost is retained. The iterative process continues until the final coefficient position is reached.
@@ III-H2 Encoder Algorithm
III-H3 Encoder Runtime Reduction
The Viterbi algorithm used in the encoder, while efficient, can still lead to a significant increase in encoder runtime. The AV2 reference encoder employs the following techniques to mitigate this complexity.
First, Single Instruction Multiple Data (SIMD) optimizations are leveraged to parallelize the trellis state evaluations. This approach yields significant reductions in encoding cycles, with a 52% decrease observed for All-Intra (AI) configurations and a 40% decrease for Random-Access (RA) configurations. Further vectorization is achieved by processing consecutive coefficients within the same diagonal coordinate in parallel, exploiting the absence of entropy context dependencies between them. Updates across trellis states are also efficiently parallelized with the execution of a small number of SIMD instructions.
Second, the complexity of the trellis search is inherently linked to the number of coefficients evaluated, which is proportional to the initial tentative EOB position. To expedite the encoding process, an EOB pre-truncation technique can be applied. This involves using a simpler, faster search method, such as rate-distortion optimized (RDO) scalar quantization, to determine a reduced EOB position. This pre-truncation effectively limits the search space for the more complex Viterbi algorithm, leading to a reduction in encoding time.
III-H4 Quantization Step Size Adjustment
For a given content and quantization step size QStep, TCQ systematically yields lower quantization distortion than its SQ counterpart. To better accommodate TCQ within the existing rate-distortion optimization (RDO) framework of AV2, the encoder adds an internal offset to q_index before deriving the value of QStep following the same mapping as in SQ. This offset is empirically chosen as +2+2 for 8-bit sequences and +4+4 for 10-bit sequences, so as to approximately align the expected distortion between SQ and TCQ at the same input q_index.
@@ III-I Probability Adaptation Rate Adjustment (PARA)
In AV2, separate adjustment parameters (denoted as $\mathbf{a}$ in Figure  10) are jointly optimized per-context (i.e., group of symbols associated with each row of the CDF table), and different time intervals defined by the $r_{C}(n)$ in Equation 13 so that each time the interval $\mathcal{T}$ is defined as:
@@ III-J1 Adaptive Coefficient Coding
ATC modifies the coefficient coding syntax in each region for AV2. In the LF region, the BR level range is extended to cover the level values $\text{BR}_{\text{LF}}=[0,1,2,3,4,\geq 5)$ using a 6-symbol alphabet. The final symbol $\geq 5$ serves as an escape character indicating that the actual value may exceed 4 which may trigger another coding pass. In the Default region, the BR range remains identical to AV1: $\text{BR}_{\text{Default}}=[0,1,2,\geq 2)$ including level values $0,1,2$ and the escape character for greater $\geq 2$.
@@
Default 2D $\min(nstats,\ 3)+\text{V}_{\text{offset}}$ 0–7
1D $\min(nstats,\ 3)+8$ 8–11
@@
For the LR coding pass the context index is derived similarly from nstats with rules varying by channel (luma or chroma) and region (LF or Default) as detailed in Table VII. Luma uses a clip of 6, while chroma uses a clip of 3. Additional offsets are applied based on coefficient position (e.g., DC vs. non-DC) in the LF region. Chroma does not have a LR pass in the LF region since BR has more symbol counts.
@@
Note: Chroma coefficients in the LF region do not use Low Range (coeff_br) coding.